OpenStudy (anonymous):

find the slope of the tangent line to the polar curve r=7/theta at theta=pi

OpenStudy (anonymous):

Use this formula \[\frac{ dy }{ dx }=\frac{ \frac{ dr }{ d \theta } }{\frac{ dr }{ d \theta } }\] with you numerator multiplied by \[\sin \theta+r \cos \theta \] and your numerator multiplied by \[\cos \theta-r \sin \theta \] After you have found you \[\frac{ dy }{ dx }\] then you will need to substitute your \[\theta=\Pi \]

OpenStudy (anonymous):

Do you get what I am trying to show you?

OpenStudy (anonymous):

If you don't go to google and search for this following link http://tutorial.math.lamar.edu/Classes/CalcII/PolarTangents.aspx

OpenStudy (anonymous):

I understand what you are saying. my only issue is what exactly do i use as x and y ?

OpenStudy (anonymous):

you need not to worry about x and y..... .in this case \[M=\frac{ dy }{ dx } slope\] You will first need to derive the given r in terms of \[\frac{ dr }{ d \theta }\] and then just plug in that in the formula that I gave you  substitute everything as it .... for example where there is r just plug it in and where there is Theta just plug it in and get your Tangent slope